L^{A}TEX News

Issue 13, June 2000

Yearly release cycle

We announced in L^{A}TEX News 11 that we intended to switch to a 12-monthly release schedule. With the present (June 2000) release, this switch is being made: thus the next release of L^{A}TEX will be dated June 2001. We shall of course continue, as in the past, to release patches as needed to fix significant bugs.

PSNFSS: Quote of the Month

You should say in the L^{A}TEX News that Walter Schmidt has taken over PSNFSS from me. It gives me a certain pleasure to be able to draw a line under that part of my life…

Sebastian Rahtz

The PSNFSS material, which supports the use of common PostScript fonts with L^{A}TEX, has been thoroughly updated. Most noticeably, the mathpple package, which used to be distributed separately, is now part of the basic PSNFSS bundle; this package provides mathematical typesetting with the Palatino typeface family. In addition, numerous bugs and flaws have been fixed and the distribution has been ‘cleaned up’. The file changes.txt contains a detailed list of these changes.

The documentation (in psnfss2e.pdf) has been completely rewritten to provide a comprehensive introduction to the use of PostScript fonts.

Notice that the new PSNFSS needs updated files for font metrics, virtual fonts and font definitions. If you received the new version (8.1) as part of a complete TEX system then these new font files should also have been installed. However, if you intend to install or update PSNFSS yourself, please read the instructions in the file 00readme.txt of the new PSNFSS distribution.

Support for commercial PostScript fonts, such as Lucida Bright, has been removed from the basic distribution; it is now available from CTAN:

http://www.tex.ac.uk/tex-archive/macros/latex/
contrib/supported/psnfssx.

New AMS-L^{A}TEX

Version 2.0 of AMS-L^{A}TEX was released on December 1, 1999. It can be obtained

via ftp://ftp.ams.org/pub/tex/ or
http://www.ams.org/tex/amslatex.html,

as well from CTAN:

http://www.tex.ac.uk/tex-archive/macros/latex/
required/amslatex.

This release consists chiefly of bug fixes and consolidation of the existing features. The division of AMS-L^{A}TEX into two main parts (the math packages; the AMS document classes) has been made more

pronounced. The files diffs-m.txt, diffs-c.txt,
amsmath.faq, and amsclass.faq describe the

changes and address some common questions.

The primary documentation files remain amsldoc.tex, for the amsmath package, and instr-l.tex, for the AMS document classes. The documentation for the amsthm package, however, has been moved from amsldoc.tex to a separate document amsthdoc.tex.

New input encoding latin4

The package inputenc has, thanks to Hana Skoumalová, been extended to cover the latin4 input encoding; this covers Baltic and Scandinavian languages as well as Greenland Inuit and Lappish.

New experimental code

In L^{A}TEX News 12 we announced some ongoing work towards a ‘Designer Interface for L^{A}TEX’ and we presented some early results thereof. Since then, at Gutenberg 2000 in Toulouse and TUG 2000 in Oxford, we described a new output routine and an improved method of handling vertical mode material between paragraphs. In combination these support higher quality automated^{1} page-breaking and page make-up for complex pages—the best yet achieved with TEX!

A paper describing the new output routine is at

http://www.latex-project.org/papers/xo-pfloat.pdf

All code examples and documentation are available at

http://www.latex-project.org/code/experimental/ .

This directory has been extended to contain

galley Prototype implementation of the interface for manipulating vertical material in galleys.

xinitials Prototype implementation of the interface for paragraph initials (needs the galley package.

xtheorem Contributed example using the template package to provide a designer interface for theorem environments.

xoutput A prototype implementation of the new output routine as described in the xo-pfloat.pdf paper. Expected availability: at or shortly after the TUG 2000 conference.

1The stress here is on automated!

L^{A}TEX News, and the L^{A}TEX software, are brought to you by the L^{A}TEX3 Project Team; Copyright 2000, all rights reserved.